# A Resolution Independent Neural Operator

### v1

@@ Abstract
The Deep operator network (DeepONet) is a powerful yet simple neural operator architecture that utilizes two deep neural networks to learn mappings between infinite-dimensional function spaces. This architecture is highly flexible, allowing the evaluation of the solution field at any location within the desired domain. However, it imposes a strict constraint on the input space, requiring all input functions to be discretized at the same locations; this limits its practical applications. In this work, we introduce a Resolution Independent Neural Operator (RINO) that provides a framework to make DeepONet resolution-independent, enabling it to handle input functions that are arbitrarily, but sufficiently finely, discretized. To this end, we propose a dictionary learning algorithm to adaptively learn a set of appropriate continuous basis functions, parameterized as implicit neural representations (INRs), from the input data. These basis functions are then used to project arbitrary input function data as a point cloud onto an embedding space (i.e., a vector space of finite dimensions) with dimensionality equal to the dictionary size, which can be directly used by DeepONet without any architectural changes. In particular, we utilize sinusoidal representation networks (SIRENs) as our trainable INR basis functions. We demonstrate the robustness and applicability of RINO in handling arbitrarily (but sufficiently richly) sampled input functions during both training and inference through several numerical examples.
@@ 1 Introduction
To achieve discretizaiton independence, we develop a data-driven algorithm that maps an arbitrarily discretized input function realization to its corresponding embedding coordinates. We construct this embedding space by projecting these input function realizations onto a set of fixed, continuous, and fully differentiable basis functions shared across all realizations, where the projection coefficients serve as the embedding coordinates. For flexibility, these basis functions are parameterized by neural networks and are learnable, although they do not have to be and could, for example, be specified a priori. Specifically, our algorithm discovers appropriate neural network basis functions from a collection of correlated signals, e.g., input function realizations. This algorithm extends standard dictionary learning techniques from finite-dimensional vector spaces to function spaces. Dictionary learning is an unsupervised machine learning method that identifies a set of basis vectors to represent data as a sparse combination of these basis vectors. It has various applications, including sparse coding [24], image denoising [19], and face recognition [37]. Unlike some classical ML methods such as principal component analysis (PCA), dictionary learning can learn overcomplete basis sets, where the number of basis vectors exceeds the dimensionality of the input space. This may provides more flexibility for signal modeling and representation.
@@
The paper is organized as follows. In Section 2, we describe the components of our proposed framework, namely the operator learning architecture, the concept of learning a dictionary of basis functions, and the parametrization approach via the INRs. In Section 3, we present four numerical examples to test the RINO framework. For each numerical example, we conducted a comprehensive study by varying the input function, the number of sensors, and their locations. Details about the data generation process are provided in each respective example. Finally, we summarize our observations and provide concluding remarks in Section 4.
@@ 2 Formulation
In Section 2.1, we revisit the operator learning problem statement that is the focus of this work. We utilize the DeepONet formalism to learn operators in a data-driven manner and develop a formulation that accepts a resolution-independent embedding of the input functions (e.g., spatially and/or temporally dependent source terms, initial conditions, material properties, etc.) instead of directly working with input function data in the original space. This approach eliminates the resolution dependency of the vanilla DeepONet from the perspective of the input function. Then, in Section 2.2, we introduce an algorithm to learn appropriate parametrized basis functions to represent input function data in a resolution-independent manner. Finally, in Section 2.3, we specifically introduce the particular parametrization of our basis function as INRs.
@@ 2.1 Operator Learning Between Function Spaces
In practice, we only have access to finite observations of input-output functions, which correspond to their discretized versions. Thus, in the most general setting, the $i$-th input function $u^{(i)}(x)$ is arbitrarily discretized at $M_{\text{in}}^{(i)}$ input sensor locations and stored as the point cloud $\mathcal{D}^{(i)}_{u}=\left\{x^{(i,j)},\bar{u}^{(i,j)}\right\}_{j=1}^{M^{(i)}_{\text{in}}}$, where $x^{(i,j)}\in\mathcal{X}$ and $\bar{u}^{(i,j)}(x^{(i,j)})\in\mathcal{U}$. Similarly, its corresponding output function $s^{(i)}(y)$ is arbitrarily discretized at $M^{(i)}_{\text{out}}$ output sensor locations and stored as a point cloud $\mathcal{D}^{(i)}_{s}=\left\{y^{(i,j)},\bar{s}^{(i,j)}\right\}_{j=1}^{M^{(i)}_{\text{out}}}$, where $y^{(i,j)}\in\mathcal{Y}$ and $\bar{s}^{(i,j)}(y^{(i,j)})\in\mathcal{S}$. This genral data configuration is schematically illustrated in Figure 1(B.1).
@@
In this manuscript, we use superscript indices to denote data-related information. Specifically, a single superscript ($\cdot^{(i)}$) indicates the $i$-th function realization, which can represent a sample. A double superscript ($\cdot^{(i,j)}$) is used to indicate the $j$-th observation point (or sensor) within the $i$-th function realization. For example, $u^{(i)}$ denotes the $i$-th realization of the function $u$, and $x^{(i,j)}$ denotes the $j$-th observation point in the $i$-th function realization. We use the bar notation $\bar{u}$ to represent the discretized form of the continuous field $u$, which may include measurement or discretization errors.
@@
$\mathcal{G}\left(u\right)(y)\approx\mathcal{F}_{\theta}(u)(y)\mathrel{\stackrel{{\scriptstyle\text{DeepONet}}}{{=}}}\sum_{k=1}^{P}\text{br}_{k}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{u}};\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (4)
where $\text{tr}_{k}(\cdot;\boldsymbol{\theta})$ and $\text{br}_{k}(\cdot;\boldsymbol{\theta})$ are parameterized functions modeled by suitable neural networks, referred to as the trunk and branch nets, respectively. The trunk net defines a set of basis functions over the domain of the output function, while the branch net defines the corresponding coefficients, which depend on the input function $u$. Within this formalism, a finite-dimensional vector representation (i.e., embedding) of the input function $u$ is necessary for use in standard neural network architectures. Indeed, the observed data $\bar{u}$ can serve as an effective representation of the input function, as used in the original DeepONet. However, this requires that the input sensor locations remain consistent across all input data, as illustrated in Figure 1(A). This has two drawbacks: first, it makes the resulting approximate operator dependent on the resolution, and second, it may cause the number of parameters in the branch network to become unnecessarily large (and therefore expensive to train), depending on how finely the input functions are discretized.
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{Q}$ is an embedding of the input function $u(x)$ in a well-defined embedding space, see Figure 1(B.2). In the next section, we will introduce a method to find this desired representation.
@@ 2.2 Dictionary Learning
In this section, we introduce a dictionary learning algorithm that identifies a set of basis functions parameterized by neural networks, which are used to approximate signals defined on point cloud data.
@@
Given a collection of $N$ signals $\boldsymbol{U}\in\mathbb{R}^{N\times M}$, which are discretized at $M$ locations and are assumed to be correlated and share a common structure, the classical dictionary learning goal is to learn a set of $Q$ basis vectors (or atoms) $\boldsymbol{\Psi}\in\mathbb{R}^{Q\times M}$ that all signals in the set can be sufficiently approximated using a sparse combination of these basis vectors (see top of Figure 2). A classical example of dictionary learning is PCA, where the goal is to approximate high-dimensional, correlated signals using a few principal components, which serve as basis vectors. Here, we go beyond discovering dictionaries that exist in finite-dimensional vector spaces and instead focus on finding those that reside in continuous function spaces, making them suitable for processing signals defined on irregular domains such as point clouds (see bottom of Figure 2).
Our goal here is to learn a dictionary that can represent the input function $u(x)$; however, the same procedure can be applied to the output function as well. In a general setup, we are given a dataset of $N$ recorded signals, denoted by $\mathcal{D}_{u}=\left\{\mathcal{D}_{u}^{(i)}\right\}_{i=1}^{N}$, where each element $\mathcal{D}_{u}^{(i)}$ contains an arbitrarily discretized version of the $i$-th function realization $u^{(i)}$, as described in the previous section and depicted in the bottom of Figure 2. If the number and locations of sensors (discretization points) are identical for all signals, the data setup resembles the classical PCA setup.
@@
These basis functions can be predefined using well-known orthogonal bases, such as Fourier bases or those used in polynomial chaos expansion (PCE) [23, 2, 12, 39, 36]. However, similar to PCA basis vectors, it would be more flexible and potentially more efficient if the basis functions could be data-dependent. To this end, we parameterize these functions as neural network functions $\psi_{l}(x;\boldsymbol{\theta}_{l})$ where $\boldsymbol{\theta}_{l}$ concatenates all parameters associated with the $l$-th neural network basis function, as follows:
@@
For the optimal values, and when $\lambda=0$, one can easily show that the residual, $r(x)=u(x)-\boldsymbol{\Psi}^{T}(x;\boldsymbol{\theta})\ \boldsymbol{\alpha}$, is orthogonal to the basis functions, hence to the subspace spanned by the dictionary. Therefore, for small values of $\lambda$ (i.e., $\lambda\to 0^{+}$), one can expect the residual to be “approximately” orthogonal to this subspace. In other words, the residual cannot be expressed linearly by the other basis functions. Therefore, it can be an effective choice as a new basis function to enhance the representation of the data. This is similar to the vector space viewpoint, where the residual vector is used as a new direction for data representation. This simple yet powerful concept has also been employed in diffusion maps (an unsupervised ML method [22, 8]) to identify unique eigendirections through local linear regression [10]. We utilize this intuitive fact to derive an algorithm that aims to iteratively construct an approximately orthogonal dictionary of atoms as descibed in Algorithm 1.
Algorithm 1 Dictionary Learning
@@
2:$\mathcal{D}_{u}=\left\{\mathcal{D}_{u}^{(i)}\right\}_{i=1}^{N},\mathcal{D}_{u}^{(i)}=\left\{\left(x^{(i,j)},\bar{u}^{(i,j)}\right)\right\}_{j=1}^{M^{(i)}}$triangleright $N$ realization of the input function $\bar{u}(x)$
@@
where $\boldsymbol{\Psi}^{(i)}(\boldsymbol{x};\boldsymbol{\theta})$ represents the basis function values at the discretization points of the $i$-th realization $\{\boldsymbol{x}^{(i,j)}\}_{j=1}^{M^{(i)}}$, and $\boldsymbol{U}^{(i)}$ are the $i$-th realization values at those corresponding points $\{\boldsymbol{u}^{(i,j)}\}_{j=1}^{M^{(i)}}$. Notice that the kernel of the linear system scales with the dictionary size and not the discretization resolution. Therefore, this implementation is usually more computationally favorable than Equation 11 which requires numerical integration and scales with the resolution size, particularly in cases where the number of basis functions is much less than the discretization resolution, which is the case in our numerical examples.

### v2

@@ Abstract
The Deep operator network (DeepONet) is a powerful yet simple neural operator architecture that utilizes two deep neural networks to learn mappings between infinite-dimensional function spaces. This architecture is highly flexible, allowing the evaluation of the solution field at any location within the desired domain. However, it imposes a strict constraint on the input space, requiring all input functions to be discretized at the same locations; this limits its practical applications. In this work, we introduce a Resolution Independent Neural Operator (RINO) that provides a framework to make DeepONet resolution-independent, enabling it to handle input functions that are arbitrarily, but sufficiently finely, discretized. To this end, we propose two dictionary learning algorithms to adaptively learn a set of appropriate continuous basis functions, parameterized as implicit neural representations (INRs), from correlated signals defined on arbitrary point cloud data. These basis functions are then used to project arbitrary input function data as a point cloud onto an embedding space (i.e., a vector space of finite dimensions) with dimensionality equal to the dictionary size, which can be directly used by DeepONet without any architectural changes. In particular, we utilize sinusoidal representation networks (SIRENs) as trainable INR basis functions. The introduced dictionary learning algorithms can be used in a similar way to learn an appropriate dictionary of basis functions for the output function data. This approach can be seen as an extension of POD-DeepONet for cases where the realizations of the output functions have different discretizations, making the Proper Orthogonal Decomposition (POD) approach inapplicable. We demonstrate the robustness and applicability of RINO in handling arbitrarily (but sufficiently richly) sampled input and output functions during both training and inference through several numerical examples.
@@ 1 Introduction
To achieve discretizaiton independence, we develop a data-driven algorithm that maps an arbitrarily discretized input function realization to its corresponding embedding coordinates. We construct this embedding space by projecting these input function realizations onto a set of fixed, continuous, and fully differentiable basis functions shared across all realizations, where the projection coefficients serve as the embedding coordinates. For flexibility, these basis functions are parameterized by neural networks and are learnable, although they do not have to be and could, for example, be specified a priori. Specifically, our algorithm discovers appropriate neural network basis functions from a collection of correlated signals, e.g., input function realizations. This algorithm extends standard dictionary learning techniques from finite-dimensional vector spaces to function spaces. Dictionary learning is an unsupervised machine learning method that identifies a set of basis vectors to represent data as a sparse combination of these basis vectors. It has various applications, including sparse coding [29], image denoising [24], and face recognition [46]. Unlike some classical ML methods such as principal component analysis (PCA), dictionary learning can learn overcomplete basis sets, where the number of basis vectors exceeds the dimensionality of the input space. This may provide more flexibility for signal modeling and representation.
@@
Connection with ReNO: The problem of interest in our study is to establish a neural operator learning method that can operate with data sampled irregularly from one realization to another while maintaining relatively consistent accuracy, particularly from the input function perspective. We assume that the function realizations are bandwidth-limited signals and that the density of sensor locations is sufficiently fine, exceeding the minimal Nyquist-Shannon sampling rate. Therefore, we do not explore potential aliasing issues related to sampling below the Nyquist-Shannon rate during training. Representation equivalent Neural Operators (ReNO) [3] were developed to investigate aliasing issues in the context of operator learning and proposes a general formalism based on frame theory to determine the conditions under which a discrete representation of an operator is equivalent to its continuous version. Since our dictionary learning algorithm identifies basis functions that are (weakly) orthogonal, they may form a frame; Bartolucci et al. [3] show that orthogonality is a sufficient condition for defining a frame. The connection between discrete and continuous representations is established by these frames (basis functions) via the synthesis operator (reconstruction operator) and its adjoint, known as the analysis operator (projection operator). Therefore, the RINO framework aligns with their frame theory, but the proposed RINO further provides an algorithm to adaptively construct appropriate basis functions (frames) to enable consistent mapping between discrete and continuous representations of signals of interest.
The paper is organized as follows. In Section 2, we describe the components of our proposed framework, namely the operator learning architecture, the concept of learning a dictionary of basis functions, and the parametrization approach via the INRs. In Section 3, we present four numerical examples to test the RINO framework. For each numerical example, we conducted a comprehensive study by varying the input function, the number of sensors, and their locations. Details about the data generation process are provided in each respective example. Finally, we summarize our observations and provide concluding remarks in Section 4. An important discussion on the connection between our proposed function-space dictionary learning algorithm and classical Gappy POD (GPOD), as well as alternative possibilities using random projections to achieve vector representations of input or output function data, is presented in Appendix F.
@@ 2 Formulation
In Section 2.1, we revisit the operator learning problem statement that is the focus of this work. We utilize the DeepONet formalism to learn operators in a data-driven manner and develop a formulation that accepts a resolution-independent embedding of the input functions (e.g., spatially and/or temporally dependent source terms, initial conditions, material properties, etc.) instead of directly working with input function data in the original space. This approach eliminates the resolution dependency of the vanilla DeepONet from the perspective of the input function. At the end of this section, we present an operator learning formulation that bypasses the need for an unknown trunk during the operator learning training. This formulation extends the POD-DeepONet approach into a fully resolution-independent framework. Then, in Section 2.2, we introduce two algorithms to learn appropriate parametrized basis functions to represent input or output function data in a resolution-independent manner. Finally, in Section 2.3, we specifically introduce the parametrization of our basis functions as INRs.
@@ 2.1 Operator Learning Between Function Spaces
In practice, we only have access to finite-dimensional observations of input-output functions, which correspond to their discretized versions. Thus, in the most general setting, the $i$-th input function $u^{(i)}(x)$ is arbitrarily discretized at $M_{\text{in}}^{(i)}$ input sensor locations and stored as the point cloud $\mathcal{D}^{(i)}_{u}=\left\{x^{(i,j)},\bar{u}^{(i,j)}\right\}_{j=1}^{M^{(i)}_{\text{in}}}$, where $x^{(i,j)}\in\mathcal{X}$ and $\bar{u}^{(i,j)}(x^{(i,j)})\in\mathcal{U}$. Similarly, its corresponding output function $s^{(i)}(y)$ is arbitrarily discretized at $M^{(i)}_{\text{out}}$ output sensor locations and stored as a point cloud $\mathcal{D}^{(i)}_{s}=\left\{y^{(i,j)},\bar{s}^{(i,j)}\right\}_{j=1}^{M^{(i)}_{\text{out}}}$, where $y^{(i,j)}\in\mathcal{Y}$ and $\bar{s}^{(i,j)}(y^{(i,j)})\in\mathcal{S}$. This general data configuration is schematically illustrated in Figure 1(B.1).
@@
In this manuscript, superscript indices are used to denote data-related information. Specifically, a single superscript ($\cdot^{(i)}$) indicates the $i$-th function realization, which can represent a sample. A double superscript ($\cdot^{(i,j)}$) is used to indicate the $j$-th observation point (or sensor) within the $i$-th function realization. For example, $u^{(i)}$ denotes the $i$-th realization of the function $u$, and $x^{(i,j)}$ denotes the $j$-th observation point in the $i$-th function realization. We use the bar notation $\bar{u}$ to represent the discretized form of the continuous field $u$, which may include measurement and/or discretization errors.
@@
$\mathcal{G}\left(u\right)(y)\approx\mathcal{F}_{\theta}(u)(y)\mathrel{\stackrel{{\scriptstyle\text{DeepONet}}}{{=}}}{\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}}^{T}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{u}})\ {\mathcal{F}_{\boldsymbol{\theta}}^{\text{tr}}}(y)=\sum_{k=1}^{P}\text{br}_{k}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bar{u}};\boldsymbol{\theta})\ \text{tr}_{k}(y;\boldsymbol{\theta}),$ (4)
where $\mathcal{F}^{\text{tr}}_{\boldsymbol{\theta}}$ and $\mathcal{F}^{\text{br}}_{\boldsymbol{\theta}}$ are parameterized functions modeled by suitable neural networks, referred to as the trunk and branch nets, respectively. The trunk net defines a set of basis functions over the domain of the output function, while the branch net defines the corresponding coefficients, which depend on the input function $u$. Within this formalism, a finite-dimensional vector representation (i.e., embedding) of the input function $u$ is necessary for use in standard neural network architectures. Indeed, the observed data $\bar{u}$ can serve as an effective representation of the input function, as used in the original DeepONet. However, this requires that the input sensor locations remain consistent across all input data, as illustrated in Figure 1(A). This has two drawbacks: first, it makes the resulting approximate operator dependent on the resolution, and second, it may cause the number of parameters in the branch network to become unnecessarily large (and therefore expensive to train), depending on how finely the input functions are discretized.
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{Q}$ is an embedding of the input function $u(x)$ in a well-defined embedding space; see Figure 1(B.2) for a schematic of the architecture and Figure 2 for a schematic of the function space mappings. Note that the selection of the embedding can be defined arbitrarily by projecting the input onto a set of continuous orthogonal basis functions (e.g. orthogonal polynomials). However, we aim to identify an optimally compact basis that both improves accuracy and reduces the number of necessary parameters in the operator. In the next section, we will introduce a method to find this desired representation.
@@
2.1.1 Operator Learning with Predefined Trunk
In the vanilla DeepONet, the trunk network essentially tries to find appropriate basis functions that approximate the output functions. So, if a suitable set of bases already exists for the output function data, it can be directly used. Using this intuition, Lu et al. [23] propose the POD-DeepONet, where the trunk network is replaced by fixed, predefined Proper Orthogonal Decomposition (POD) modes associated with the output functions, resulting in only the branch network being trainable. This approach may facilitate training and improve the accuracy of the operator learning task. However, due to the use of POD, which works on fixed grids, similar discretization dependence issues arise, as with the input function process in the vanilla DeepONet. For example, if the discretization of output functions varies from realization to realization, POD cannot be directly used.
In RINO, we develop dictionary learning algorithms that adaptively learn basis functions for approximating signals defined on arbitrary point clouds. This allows us to project arbitrary signals onto these basis functions and use their corresponding coefficients as appropriate representations in the embedding space. This is the nature of the embedding of the input function in Eq. (5). If such a method exists (as discussed in the next section), it can also be applied separately to the output function data. This process would build an appropriate dictionary of basis functions for the output data, which can subsequently be used in the operator learning task as a fixed, predefined trunk (analogous to the POD) as showed in Figure 2(b).
Figure 2: Operator learning with RINO for (a) unknown trunk and (b) predefined trunk basis functions. In (a), only the input functions are embedded using a learned set of basis function and the operator is learned between the embedding space and the output directly by training both branch ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) and trunk ($\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$) networks. In (b), is functions are learned separately for both output and input function data. Consequently, the operator learning task effectively becomes a mapping between these two constructed subspaces via the branch network $\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}$.
For now, let us additionally assume that the output dictionary $\boldsymbol{\Psi}_{S}(y)=\{\psi_{l}^{S}(y)\}_{l=1}^{P}$ is known and available for use where $P=|\Psi_{S}|$. The availability of the input dictionary, denoted by $\boldsymbol{\Psi}_{U}(x)=\{\psi_{l}^{U}(x)\}_{l=1}^{Q}$ where $Q=|\Psi_{U}|$, was already assumed in the previous section. Under this setup, which is schematically shown in Figure 2(b), the objective function for operator learning is defined as:
$\underset{\boldsymbol{\theta}}{\text{argmin}}\ \mathbb{E}_{i}\mathbb{E}_{(y,s)\sim\mathcal{D}_{s}^{(i)}}\underbrace{\left\|s-\text{Recon}_{\Psi_{S}}\left(y;\bar{\boldsymbol{\gamma}}^{(i)}_{\boldsymbol{\theta}}\right)\right\|_{2}^{2}}_{\text{prediction loss}}+\tau\underbrace{\left\|\boldsymbol{\gamma}^{(i)}-\bar{\boldsymbol{\gamma}}^{(i)}_{\boldsymbol{\theta}}\right\|_{2}^{2}}_{\text{embedding consistency loss}},$ (7)
where $\bar{\boldsymbol{\gamma}}^{(i)}_{\boldsymbol{\theta}}=\mathcal{F}_{\boldsymbol{\theta}}^{\text{br}}(\boldsymbol{\alpha}^{(i)})$, $\boldsymbol{\gamma}^{(i)}=\text{Proj}_{\Psi_{S}}(\mathcal{D}^{(i)}_{s})$, $\boldsymbol{\alpha}^{(i)}=\text{Proj}_{\Psi_{U}}(\mathcal{D}^{(i)}_{u})$, and $\tau$ is a penalty factor that controls the extent to which the second term is enforced. The first term in this objective function corresponds to the prediction part of the operator network, while the second term helps the branch network produce a representation consistent with the coefficients of the output basis functions. The reconstruction operator for a dictionary $\boldsymbol{\Psi}$ is defined as follows:
$\text{Recon}_{\Psi}(\cdot;\boldsymbol{\gamma})\triangleq\sum_{l=1}^{|\Psi|}\ \psi_{l}(\cdot)\ \gamma_{l}.$ (8)
The dictionary projection operator will be specified in the next section.
@@ 2.2 Dictionary Learning
In this section, we introduce two dictionary learning algorithms to identify a set of basis functions parameterized by neural networks, which are used to approximate signals defined on point cloud data. The first algorithm aims to adaptively find basis functions that capture the expected value of the residual across batches of signal realizations, while the second algorithm aims to find a basis each time that is needed to approximate only one realization at a time.
2.2.1 Batch-wise Learning
Given a collection of $N$ signals $\boldsymbol{U}\in\mathbb{R}^{N\times M}$, which are discretized at $M$ locations and assumed to be correlated and share a common structure, the classical dictionary learning goal is to learn a set of $Q$ basis vectors (or atoms), $\boldsymbol{\Psi}\in\mathbb{R}^{Q\times M}$, such that all signals in the set can be sufficiently approximated using a sparse linear combination of these basis vectors (see top of Figure 3). A classical example of dictionary learning is PCA, where the goal is to approximate high-dimensional, correlated signals using a few principal components, which serve as basis vectors. Here, we go beyond discovering dictionaries that exist in finite-dimensional vector spaces and instead focus on finding those that reside in continuous function spaces, making them suitable for processing signals defined on irregular domains such as point clouds (see bottom of Figure 3).
@@ 2.2.1 Batch-wise Learning
Our goal here is to learn a dictionary that can represent the input function $u(x)$; however, the same procedure can be applied to the output function as well. In a general setup, we are given a dataset of $N$ recorded signals, denoted by $\mathcal{D}_{u}=\{\mathcal{D}_{u}^{(i)}\}_{i=1}^{N}$, where each element $\mathcal{D}_{u}^{(i)}$ contains an arbitrarily discretized version of the $i$-th function realization $u^{(i)}$, as described in the previous section and depicted in the bottom of Figure 3. If the number and locations of sensors (discretization points) are identical for all signals, the data setup resembles the classical PCA setup.
@@
These basis functions can be predefined using well-known orthogonal bases, such as Fourier bases or those used in polynomial chaos expansion (PCE) [28, 4, 15, 48, 45]. However, similar to PCA basis vectors, it would be more flexible and potentially more efficient if the basis functions could be data-dependent. To this end, we parameterize these functions as neural networks $\psi_{l}(x;\boldsymbol{\theta}_{l})$ where $\boldsymbol{\theta}_{l}$ concatenates all parameters associated with the $l$-th neural network basis function, as follows:
@@
For the optimal values, and when $\lambda=0$, one can easily show that the residual, $r(x)=u(x)-\boldsymbol{\Psi}^{T}(x;\boldsymbol{\theta})\ \boldsymbol{\alpha}$, is orthogonal to the basis functions, hence to the subspace spanned by the dictionary. Therefore, for small values of $\lambda$ (i.e., $\lambda\to 0^{+}$), one can expect the residual to be “approximately” orthogonal to this subspace. In other words, the residual cannot be expressed linearly by the other basis functions. Therefore, the residual can serve as an effective choice for a new basis function to enhance the representation of the data, while maintaining (approximate) orthogonality. This is similar to the vector space viewpoint, where the residual vector is used as a new direction for data representation. This simple yet powerful concept has also been employed in diffusion maps (an unsupervised ML method [27, 10]) to identify unique eigendirections through local linear regression [12]. We utilize this intuitive fact to derive an algorithm that aims to iteratively construct an approximately orthogonal dictionary of atoms as descibed in Algorithm 1.
Algorithm 1 Dictionary Learning (Batch-wise)
@@
2:$\mathcal{D}_{u}=\left\{\mathcal{D}_{u}^{(i)}\right\}_{i=1}^{N},\mathcal{D}_{u}^{(i)}=\left\{\left(x^{(i,j)},\bar{u}^{(i,j)}\right)\right\}_{j=1}^{M^{(i)}}$triangleright $N$ realizations of the input function $\bar{u}(x)$
@@
where $\boldsymbol{\Psi}^{(i)}(\boldsymbol{x};\boldsymbol{\theta})$ represents the basis function values at the discretization points of the $i$-th realization $\{\boldsymbol{x}^{(i,j)}\}_{j=1}^{M^{(i)}}$, and $\boldsymbol{U}^{(i)}$ are the $i$-th realization values at those corresponding points $\{\boldsymbol{u}^{(i,j)}\}_{j=1}^{M^{(i)}}$. Notice that the kernel of the linear system scales with the dictionary size and not the discretization resolution. Therefore, this implementation is usually more computationally favorable than Eq. (13) which requires numerical integration and scales with the resolution size, particularly in cases where the number of basis functions is much less than the discretization resolution, which is the case in our numerical examples.